# Realising VCD for untwisted automorphism groups of RAAGs

### v1

@@ 4.5. Resultant complex
Conversely, the only way for a cube to be retracted is to either contain hugged partitions, or be the face along which we push in. In the latter case, necessarily we can add a non-principal partition $\mathcal{Q}$ to $\Pi_{2}$ which becomes hugged in $\Pi_{2}\cup\mathcal{Q}$. ∎
@@
Observe that as the $U(A_{\Gamma})$-action on $\widehat{K_{\Gamma}}$ is proper and cocompact, we know that the virtual cohomological dimension of $U(A_{\Gamma})$ satisfies
@@ 5. Realising virtual cohomological dimension
Millard–Vogtmann [MV19] prove that if $\Gamma$ is ‘barbed’ and $M(V)>M(L)$, then every top-dimensional cube in $K_{\Gamma}$ has a free face; this allows one to perform an equivariant retraction of $K_{\Gamma}$ which reduces its dimension by one. In this section, we strengthen this result as follows: if $\Gamma$ is ‘barbed’ then our resultant complex $\widehat{K_{\Gamma}}$ always has dimension exactly $M(L)$.
We use the following definition from [MV19].
@@
Let $\Gamma$ be barbed. Then any oversize set of $\Gamma$-Whitehead partitions contains a hugged partition.
@@
Let $\Pi$ be an oversize compatible set of $\Gamma$-partitions. Then $\Pi$ contains at least one non-principal partition; let $\mathcal{Q}$ be a non-principal partition such that one of its sides (say, $Q$) contains no sides of any other non-principal partitions.
@@
Now suppose that there is some ′′$\mathcal{R}^{\prime\prime}\in\Pi$ with which $\mathcal{P}$ is incompatible. Let $\mathcal{R}^{\prime\prime}$ be based at $n$; by what we just said, we must have $n\neq m$. Since $\mathcal{P}$ and $\mathcal{R}^{\prime\prime}$ are incompatible, we must have $d_{\Gamma}(m,n)=2$. Since $m>_{\circ}u$, we must also have $d_{\Gamma}(n,u)\geq 2$ (as otherwise $[n,u]=1$ which implies that $[m,n]=1$).
@@
This process must terminate at some point, since $Q$ contains only finitely many vertices. If it terminates by finding that $\mathcal{Q}$ is hugged, then we’re done, so suppose that is not the case. This means that it terminated because we found some principal partition $\mathcal{P}$ which could replace $\mathcal{Q}$, to form $\Pi^{\prime}\coloneqq\left(\Pi\setminus\mathcal{Q})\cup\mathcal{P}\right)$.
We now go back to the start of Step 1 with this new $\Pi^{\prime}$: pick a non-principal partition with one side containing no sides of any other non-principal partitions, show that it is hugged in $\Pi^{\prime}$, or replace it too by a principal partition. This process must terminate, as otherwise we will successfully replace all the non-principal partitions in $\Pi$ by principal partitions, resulting in an oversize set of pairwise compatible principal partitions—a contradiction of the definition of ‘oversized’.
@@
Claim: $\widehat{\mathcal{Q}}$ is 2-hugged in $\widehat{\Pi}$.
Suppose that $\widehat{\mathcal{Q}}$ is 1-hugged in $\widehat{\Pi}$ by some principal partition $\mathcal{P}$; let $\widehat{Q}$ be the side of $\widehat{\mathcal{Q}}$ which is hugged. If $\mathcal{P}\in\Pi$, we’re done; if not, then $\mathcal{P}$ was added in at some point, in replacement of some non-principal partition $\mathcal{Q}$ (which is based at, say, $u$). Let $\mathcal{P}$ be based at some principal vertex $m$; it had a side $P\subseteq Q$. We have $m>_{\circ}u$ and $m>_{\circ}\widehat{u}$, so we cannot have $d_{\Gamma}(u,\widehat{u})=1$. Since $\mathcal{Q}$ was chosen so that $Q$ was innermost among non-principal sides, we cannot have a $\widehat{Q}$ hugged by the side $P$—this would mean that $\widehat{Q}\subseteq Q$. Hence $\widehat{Q}$ must be hugged by the other side $\overline{P}$ of $\mathcal{P}$. Therefore $\widehat{Q}\supseteq\{m,m^{-1}\}$, which (since $d_{\Gamma}(u,\widehat{u})\neq 1$) means that $\widehat{Q}\supseteq Q$. But now $P\cup\overline{P}\subseteq\widehat{Q}$, which implies that $\widehat{\mathcal{Q}}$ is not a valid partition: one of its sides is a singleton. We conclude that $\widehat{\mathcal{Q}}$ is not 1-hugged in $\widehat{\Pi}$.
This proves the claim; suppose that $\widehat{\mathcal{Q}}$ is 2-hugged in $\widehat{\Pi}$ by $\{\mathcal{P},\mathcal{P}^{\prime}\}$.
@@
[…]
By compatibility of $\mathcal{Q}$ with $\mathcal{P}^{\prime}$, and since $Q$ contains $\{m,m^{-1}\}$, we know that $Q$ must contain a side of $\mathcal{P}^{\prime}$. If $P^{\prime}\subseteq Q$, then $\smash{\overline{\widehat{Q}}=\overline{P}\cap\overline{P^{\prime}}\supseteq\overline{Q}}$, which implies that $\widehat{Q}\subseteq Q$. Hence (again as $Q$ was chosen to be innermost among non-principal sides in a set that contained $\widehat{\mathcal{Q}}$) we have $Q=Q^{\prime}$, a contradiction (as $\widehat{\mathcal{Q}}\in\widehat{\Pi}$, a set that does not contain $\mathcal{Q}$).
@@
If $\Gamma$ is spiky, then every cube in $K_{\Gamma}$ of dimension greater than $M(L)$ is redundant.
@@
Since $\Gamma$ is barbed, we may apply the retraction process to $K_{\Gamma}$, obtaining the new cube complex $\widehat{K_{\Gamma}}$.
A cube $c(\Pi_{1},\Pi_{2};\alpha)$ in $K_{\Gamma}$ will have dimension strictly greater than $M(L)$ if and only if $\Pi_{2}\setminus\Pi_{1}$ is an oversize set. By Lemma 5.4, and since $\Gamma$ is barbed, we know that $\Pi_{2}\setminus\Pi_{1}$ contains a non-principal partition which is hugged in $\Pi_{2}\setminus\Pi_{1}$—so in particular, is hugged in $\Pi_{2}$. Now by the characterisation of $\widehat{K_{\Gamma}}$ in Proposition 4.19, all such cubes were removed during the retraction process and so are not present in $\widehat{K_{\Gamma}}$.
@@ 6. Application: Arbitrarily large gaps between $\textsc{vcd}(U(A_{\Gamma}))$ and $\dim(K_{\Gamma})$
In this section we present a certain prototypical family of examples, which we call rake graphs. We explicitly calculate $M(L)$ and $M(V)$ for these graphs; we then apply Theorem 5.6 to show that $\textsc{vcd}(U(A_{\Gamma}))=M(L)$ and hence that the gap between $\textsc{vcd}(U(A_{\Gamma}))$ and $\dim(K_{\Gamma})$ may be arbitrarily large.
@@ 6.1. Rake graphs
Since $T_{d}^{\pm}\setminus\operatorname{st}({v})^{\pm}$ consists only of the $b_{i}^{\pm}$ and no edges, we have $M(v)=2d-1$. Indeed, the $v$-sides of a compatible set of partitions all based at $v$ are totally ordered by inclusion. The smallest possible $v$-side consists of two vertices: $v$ and one of $\{b_{1},\overline{b_{1}},\dots,b_{d},\overline{b_{d}}\}$; the largest possible consists of $2d$ vertices: $v$ along with all but one of $\{b_{1},\overline{b_{1}},\dots,b_{d},\overline{b_{d}}\}$. Hence $M(v)\leq 2d-1$. Equality can be realised, for example by the following set of $T_{d}$-partitions (given by their $v$-sides):
@@
We now prove that $M(a_{1},\dots,a_{d})=d$. Note that partitions based at $a_{i}$ have $|\mathcal{C}(a_{i})|=d+1$: we have the components $\{u\},\{\overline{u}\}$, and $S_{j}\coloneqq\{a_{j}^{\pm},b_{j}^{\pm}\}$ for $j\neq i$. Hence $M(a_{i})\leq d$ for each $i$. Let $\mathcal{P}_{i}$ be based at $a_{i}$ and $\mathcal{P}_{j}$ based at $a_{j}$$i\neq j$. Then, since $[a_{i},a_{j}]\neq 1$, we must have a choice of sides $P_{i}^{\times},P_{j}^{\times}$ such that $P_{i}^{\times}\cap P_{j}^{\times}=\emptyset$$\mathcal{P}_{i}$ splits $a_{i}$, so $P_{j}^{\times}$ must not contain the connected component $S_{i}\in\mathcal{C}(a_{j})$. Hence $\overline{P_{j}^{\times}}$ contains this connected component. Similarly, $P_{i}^{\times}$ does not contain the connected component $S_{j}\in\mathcal{C}(a_{i})$. Now, every element of $\mathcal{C}(a_{i})$ which is contained in $P_{i}^{\times}$ is contained in $\overline{P_{j}^{\times}}$, as otherwise we could not have $P_{i}^{\times}\cap P_{j}^{\times}=\emptyset$. Hence $P_{i}^{\times}\subsetneq\overline{P_{j}^{\times}}$. Moreover, $\overline{P_{j}^{\times}}$ contains at least one more element of $\mathcal{C}(a_{j})$ than the number of elements of $\mathcal{C}(a_{i})$ contained in $P_{i}^{\times}$. Hence, any set of pairwise compatible partitions all based at elements of $\{a_{1},\dots,a_{d}\}$ form a chain of sides totally ordered by inclusion, and this chain can be of length at most $d$, so $M(a_{1},\dots,a_{d})\leq d$. One can reapply this argument to show that $M(W)\leq d$, for any subset $W\subseteq\{a_{1},\dots,a_{d}\}$. Equality can also be realised here. For example, the following is a set of size $d$ of partitions, one based at $a_{i}$ for each $i=1,\dots,d$:
[…]
Since $M(L)=M(v)+M(a_{1},\dots,a_{d})$, we now have $M(L)=2d-1+d=3d-1$.
The only non-principal vertex at which partitions can be based is $u$. Any set of pairwise compatible partitions, all based at $u$, must have their $u$-sides totally ordered by inclusion. The only elements of $\mathcal{C}(u)$ are $S_{1},\dots,S_{d}$; since there are $d$ of these we therefore have $M(u)=d-1$, using the same reasoning as for $v$ above.
@@
[…]
@@ 6.2. Arbitrarily large gaps
Applying the work of Millard–Vogtmann [MV19] to the calculations from the previous subsection, we have the bounds
$23d-1=M(L)\leq\textsc{vcd}(U(A_{T_{d}}))\leq M(V)=\dim(K_{T_{d}})=4d-$ 2 .
In fact, the rake graphs are all barbed, so ([MV19], Theorem 7.8) yields the tighter upper bound of $34d-$ 3. In particular, we have $\textsc{vcd}(U(A_{T_{2}}))=5$.
@@
where $e$ is the number of edges and $\ell$ is the number of leaves. As noted in ([CSV17], §5.2), since the normal subgroup of twists $T(A_{\Gamma})\leq\operatorname{Out}(A_{\Gamma})$ is free abelian of rank $\ell$ and since the intersection $T(A_{\Gamma})\cap U(A_{\Gamma})=\{1\}$, it is natural to hope that $\textsc{vcd}(U(A_{\Gamma}))=e+\ell-3$. For the $d$-rake, this number is $(2d+1)+(d+1)-3=3d-1=M(L)$.
@@
    the difference between $\dim(\widehat{K_{\Gamma}})$ and $\dim(K_{\Gamma})$ is arbitrarily large;
@@ 7.1. A sufficient condition for $M(L)=M(V)$
($\dagger$) all principal maximal m>u are in the same component of $\text{all principal maximal~{}$m>_{\circ}u$ are in the same component of~{}$\Gamma\setminus\operatorname{st}({u})$}.$
@@
[…]
@@
[…]
@@
Consider any compatible set $\Pi$ of partitions based in the set $\{a_{1},a_{3}\}$. We consider how many partitions based at $a_{2}$ it is possible to add to $\Pi$ to obtain a set of size $M(a_{1},a_{2},a_{3})$. If $\Pi$ is empty, we can add at most five partitions based at $a_{2}$, since $M(a_{2})=5$. Assume $\Pi\neq\emptyset$; without loss of generality let $\mathcal{P}_{1}\in\Pi$ be based at $a_{1}$. We know that this has one side containing one or both of the components $\{u_{1}\}$$\{\overline{u_{1}}\}$. But these are both elements of $\mathcal{C}(a_{2})$, so if there is a partition based at $a_{1}$ with both of these components on the same side, then we can only add at most three partitions based at $a_{2}$. Similarly, if there is no partition based at $a_{1}$ with both of these components on the same side, then we may add at most four partitions based at $a_{2}$. Now a symmetric argument with any partitions based at $a_{3}$ further reduces the options for partitions based at $a_{2}$ which can be added to $\Pi$ to obtain a larger compatible set. If there is a partition based at $a_{3}$ with both of the components $\{u_{3}\}$$\{\overline{u_{3}}\}$ on the same side, then the number of partitions based at $a_{2}$ we can add is reduced by another two. If there is only one partition based at $a_{3}$ and the components $\{u_{3}\}$ and $\{\overline{u_{3}}\}$ are on different sides, then the number of partitions based at $a_{2}$ which we can add is further reduced by only one. Notice that if $\Pi$ contains two partitions based at $a_{1}$ then one of them necessarily has $\{u_{1}\}$ and $\{\overline{u_{1}}\}$ on the same side; the same holds for $a_{3}$. Hence $M(a_{1},a_{2},a_{3})\leq 5$ (and since $M(a_{2})=5$, we have $M(a_{1},a_{2},a_{3})=5$).
Hence $M(L)\leq M(v_{1},v_{2})+M(a_{1},a_{2},a_{3})=6+5=11$. In fact we have equality, as witnessed by the following compatible set of $\Delta$-partitions, defined by the following sides:
@@
We would like to apply our work to show that $\textsc{vcd}(U(A_{\Delta}))=M(L)$. However, $\Delta$ violates Condition 2, so we cannot use Lemma 4.13. It turns out that $\Delta$ actually still satisfies the conclusion of Lemma 4.13, which we now prove by hand. Since our work only relies on graphs satisfying Condition 2 so that Lemma 4.13 is applicable, this is sufficient to apply Theorem 5.6.
Let $\Pi$ be a set of compatible $\Delta$-partitions. Suppose that $\Pi$ contains non-principal partitions $\mathcal{Q}$$\mathcal{Q}^{\prime}$ which are hugged in $\Pi$. We would like to show that if $\mathcal{R}$ is some principal partition which is compatible with all the principal partitions in $\Pi$ (so in particular, with all partitions involved in hugging $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$), then $\mathcal{R}$ is compatible with at least one of $\{\mathcal{Q},\mathcal{Q}^{\prime}\}$.

### v2

@@ 4.5. Resultant complex
Conversely, the only way for a cube $c(\Pi_{1},\Pi_{2})$ to be retracted is for either $\Pi_{2}\setminus\Pi_{1}$ to contain hugged partitions, or to be the face along which we push in. In the latter case, necessarily we can add a non-principal partition $\mathcal{Q}$ to $\Pi_{2}$ which becomes hugged in $\Pi_{2}\cup\mathcal{Q}$. ∎
@@
Remark 4.23.
Note that it is possible for single vertices in $K_{\Gamma}$ to be redundant. For example, take $\Pi_{1}=\{\mathcal{P}_{1},\mathcal{P}_{2}\}$ and $\Pi_{2}=\{\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{Q}\}$, where $\mathcal{Q}$ is hugged by $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Then it may happen that at some stage, the (1-dimensional) cube $c(\Pi_{1},\Pi_{2})$ (which is redundant since $\Pi_{2}\setminus\Pi_{1}=\mathcal{Q}$ is hugged in $\Pi_{2}$) gets removed via its (0-dimensional) subface $c(\Pi_{1},\Pi_{1})$ (which is redundant since $\mathcal{Q}$ can be added to $\Pi_{1}$ to become hugged in $\Pi_{1}\cup\mathcal{Q}=\Pi_{2}$). Hence the vertex sets of $K_{\Gamma}$ and $\widehat{K_{\Gamma}}$ may not coincide.
Observe (applying Theorem 2.7) that as the $U(A_{\Gamma})$-action on $\widehat{K_{\Gamma}}$ is proper and cocompact, we know that the virtual cohomological dimension of $U(A_{\Gamma})$ satisfies
@@ 5. Realising virtual cohomological dimension
Millard–Vogtmann [MV21] prove that if $\Gamma$ is ‘barbed’ and $M(V)>M(L)$, then every top-dimensional cube in $K_{\Gamma}$ has a free face; this allows one to perform an equivariant retraction of $K_{\Gamma}$ which reduces its dimension by one. In this section, we strengthen this result as follows: if $\Gamma$ is ‘barbed’ then our resultant complex $\widehat{K_{\Gamma}}$ always has dimension exactly $M(L)$.
We use the following definition from [MV21].
@@
Let $\Gamma$ be barbed. Then any oversize set of compatible $\Gamma$-Whitehead partitions contains a hugged partition.
@@
Let $\Pi$ be an oversize set of compatible $\Gamma$-partitions. Then $\Pi$ contains at least one non-principal partition; let $\mathcal{Q}$ be a non-principal partition such that one of its sides (say, $Q$) contains no sides of any other non-principal partitions.
@@
Now suppose that there is some $\mathcal{R}^{\prime\prime}\in\Pi$ with which $\mathcal{P}$ is incompatible. Let $\mathcal{R}^{\prime\prime}$ be based at $n$; by what we just said, we must have $n\neq m$. Since $\mathcal{P}$ and $\mathcal{R}^{\prime\prime}$ are incompatible, we must have $d_{\Gamma}(m,n)=2$. Since $m>_{\circ}u$, we must also have $d_{\Gamma}(n,u)\geq 2$ (as otherwise $[n,u]=1$ which implies that $[m,n]=1$).
@@
This process must terminate at some point, since $Q$ contains only finitely many vertices. If it terminates by finding that $\mathcal{Q}$ is hugged, then we’re done, so suppose that is not the case. This means that it terminated because we found some principal partition $\mathcal{P}$ which could replace $\mathcal{Q}$, to form $\Pi^{\prime}\coloneqq\left(\Pi\setminus\mathcal{Q}\right)\cup\mathcal{P}$.
We now go back to the start of Step 1 with this new $\Pi^{\prime}$: pick a non-principal partition with one side containing no sides of any other non-principal partitions, show that it is hugged in $\Pi^{\prime}$, or replace it too by a principal partition. This process must terminate, as otherwise we will successfully replace all the non-principal partitions in $\Pi$ by principal partitions, resulting in an oversize set of pairwise compatible principal partitions – a contradiction of the definition of ‘oversized’.
@@
Claim.
$\widehat{\mathcal{Q}}$ is 2-hugged in $\widehat{\Pi}$.
Proof of Claim..
[…]
We can now suppose that $\widehat{\mathcal{Q}}$ is 2-hugged in $\widehat{\Pi}$ by $\{\mathcal{P},\mathcal{P}^{\prime}\}$.
@@
Since $d_{\Gamma}(u,u^{\prime})\neq 1$, we know that $Q$ either contains, or is contained in, one side of $\widehat{\mathcal{Q}}$. We chose $Q$ to be innermost among non-principal sides of a set of partitions that contained $\widehat{\mathcal{Q}}$, so we must be in the latter case. Since $m\in Q\cap\widehat{Q}$, we therefore conclude that $Q\subseteq\widehat{Q}$.
By compatibility of $\mathcal{Q}$ with $\mathcal{P}^{\prime}$, and since $Q$ contains $\{m,m^{-1}\}$, we know that $Q$ must contain a side of $\mathcal{P}^{\prime}$. If $P^{\prime}\subseteq Q$, then $\smash{\overline{\widehat{Q}}=\overline{P}\cap\overline{P^{\prime}}\supseteq\overline{Q}}$, which implies that $\widehat{Q}\subseteq Q$. Hence (again as $Q$ was chosen to be innermost among non-principal sides of a set of partitions that contained $\widehat{\mathcal{Q}}$) we have $Q=Q^{\prime}$, a contradiction (as $\widehat{\mathcal{Q}}\in\widehat{\Pi}$, a set that does not contain $\mathcal{Q}$).
@@
If $\Gamma$ is barbed, then every cube in $K_{\Gamma}$ of dimension greater than the principal rank $M(L)$ is redundant.
@@
Since $\Gamma$ is spiky, we may apply the retraction process to $K_{\Gamma}$, obtaining the new $U(A_{\Gamma})$-complex $\widehat{K_{\Gamma}}$.
A cube $c(\Pi_{1},\Pi_{2};\alpha)$ in $K_{\Gamma}$ will have dimension strictly greater than $M(L)$ if and only if $\Pi_{2}\setminus\Pi_{1}$ is an oversize set. By Lemma 5.4, and since $\Gamma$ is barbed, we know that $\Pi_{2}\setminus\Pi_{1}$ contains a non-principal partition which is hugged in $\Pi_{2}\setminus\Pi_{1}$ – so in particular, is hugged in $\Pi_{2}$. Now by the characterisation of $\widehat{K_{\Gamma}}$ in Proposition 4.20, all such cubes were removed during the retraction process and so are not present in $\widehat{K_{\Gamma}}$.
@@ 6. Application: Arbitrarily large gaps between $\textsc{vcd}(U(A_{\Gamma}))$ and $\dim(K_{\Gamma})$
In this section we present a certain prototypical family of examples, which we call rake graphs. We explicitly calculate $M(L)$ and $M(V)$ for these graphs; we then apply Corollary 5.6 to show that $\textsc{vcd}(U(A_{\Gamma}))=M(L)$ and hence that the gap between $\textsc{vcd}(U(A_{\Gamma}))$ and $\dim(K_{\Gamma})$ may be arbitrarily large.
@@ 6.1. Rake graphs
Since $T_{d}^{\pm}\setminus\operatorname{st}({v})^{\pm}$ consists only of the $b_{i}^{\pm}$ and no edges, we have M(v)=2d1M(v)=2d-1. Indeed, the $v$-sides of a compatible set of partitions all based at $v$ are totally ordered by inclusion. The smallest possible $v$-side consists of two vertices: $v$ and one of $\{b_{1},\overline{b_{1}},\dots,b_{d},\overline{b_{d}}\}$; the largest possible consists of $2d$ vertices: $v$ along with all but one of $\{b_{1},\overline{b_{1}},\dots,b_{d},\overline{b_{d}}\}$. Hence $M(v)\leq 2d-1$. Equality can be realised, for example by the following set of $T_{d}$-partitions (given by their $v$-sides):
@@
We now prove that $M(a_{1},\dots,a_{d})=d$. Note that partitions based at $a_{i}$ have $|\mathcal{C}(a_{i})|=d+1$: we have the components $\{u\},\{\overline{u}\}$, and $S_{j}\coloneqq\{a_{j}^{\pm},b_{j}^{\pm}\}$ for $j\neq i$. Hence $M(a_{i})\leq d$ for each $i$. Let $\mathcal{P}_{i}$ be based at $a_{i}$ and $\mathcal{P}_{j}$ based at $a_{j}$$i\neq j$. Then, since $a_{i},a_{j}]\neq 1$, we must have a choice of sides $P_{i}^{\times},P_{j}^{\times}$ such that $P_{i}^{\times}\cap P_{j}^{\times}=\emptyset$$\mathcal{P}_{i}$ splits $a_{i}$, so $P_{j}^{\times}$ must not contain the connected component $S_{i}\in\mathcal{C}(a_{j})$. Hence $\overline{P_{j}^{\times}}$ contains this connected component. Similarly, $P_{i}^{\times}$ does not contain the connected component $S_{j}\in\mathcal{C}(a_{i})$. Now, every element of $\mathcal{C}(a_{i})$ which is contained in $P_{i}^{\times}$ is contained in $\overline{P_{j}^{\times}}$, as otherwise we could not have $P_{i}^{\times}\cap P_{j}^{\times}=\emptyset$. Hence $P_{i}^{\times}\subsetneq\overline{P_{j}^{\times}}$. Moreover, $\overline{P_{j}^{\times}}$ contains at least one more element of $\mathcal{C}(a_{j})$ than the number of elements of $\mathcal{C}(a_{i})$ contained in $P_{i}^{\times}$. Hence, any set of pairwise compatible partitions all based at elements of $\{a_{1},\dots,a_{d}\}$ form a chain of sides totally ordered by inclusion, and this chain can be of length at most $d$, so $M(a_{1},\dots,a_{d})\leq d$. One can reapply this argument to show that $M(W)\leq d$, for any subset $W\subseteq\{a_{1},\dots,a_{d}\}$. Equality can also be realised here. For example, the following is a set of size $d$ of partitions, one based at $a_{i}$ for each $i=1,\dots,d$:
($\star\star$) $\{a_{1},u\},\;\;\{a_{2},u,S_{1}\},\;\;\dots\;\;,\{a_{d},u,S_{1},\dots,S_{d-1}\}.$
Since $M(L)=M(v)+M(a_{1},\dots,a_{d})$, we now have M(L)=2d1+d=3d1M(L)=2d-1+d=3d-1, as required.
The only non-principal vertex at which partitions can be based is $u$. Any set of pairwise compatible partitions, all based at $u$, must have their $u$-sides totally ordered by inclusion. The only elements of $\mathcal{C}(u)$ are $S_{1},\dots,S_{d}$; since there are $d$ of these we therefore have M(u)=d1M(u)=d-1, using the same reasoning as for $v$ above.
@@
together with the sets given in ($\star$6.1) and ($\star\star$6.1) gives a set of pairwise compatible $T_{d}$-partitions which realises the equality M(V)=4d2M(V)=4d-2. ∎
@@ 6.2. Arbitrarily large gaps
Applying the work of Millard–Vogtmann [MV21] to the calculations from the previous subsection, we have the bounds
$3d-1=M(L)\leq\textsc{vcd}(U(A_{T_{d}}))\leq M(V)=\dim(K_{T_{d}})=4d-2.$
In fact, the rake graphs are all barbed, so ([MV21], Theorem 7.8) yields the tighter upper bound of 4d34d-3. In particular, we have $\textsc{vcd}(U(A_{T_{2}}))=5$.
@@
where $e$ is the number of edges and $\ell$ is the number of leaves. As noted in ([CSV17], §5.2), since the normal subgroup of twists $T(A_{\Gamma})\leq\operatorname{Out}(A_{\Gamma})$ is free abelian of rank $\ell$ and since the intersection $T(A_{\Gamma})\cap U(A_{\Gamma})=\{1\}$, it is natural to hope that $\textsc{vcd}(U(A_{\Gamma}))=e+\ell-3$. For the $d$-rake, this number is (2d+1)+(d+1)3=3d1=M(L)(2d+1)+(d+1)-3=3d-1=M(L).
@@
[…]
@@ 7.1. A sufficient condition for $M(L)=M(V)$
[…]
@@
The $d$-rake violates ($\dagger$7.1) in a rather controlled manner, and the gap between $\textsc{vcd}(U(A_{T_{d}}))$ and $\dim(K_{T_{d}})$ is predictable: there are precisely $d$ components of $T_{d}\setminus\operatorname{st}({u})$ containing a principal maximal $m>_{\circ}u$, and the gap between $\textsc{vcd}(U(A_{T_{d}}))=M(L)$ and $\dim(K_{T_{d}})=M(V)$ is precisely d1d-1. One might therefore consider a statement along the lines of:
@@
The elements of $\mathcal{C}(v_{1})$ are $\{b_{1}\},\{\overline{b_{1}}\},\{b_{2}\},\{\overline{b_{2}}\},C\coloneqq\{u_{2}^{\pm},v_{2}^{\pm},a_{3}^{\pm},b_{3}^{\pm}\}$. Since there are five components here, we have $M(v_{1})\leq 4$ (and we can realise equality here). The elements of $\mathcal{C}(v_{2})$ are similar, and $M(v_{2})=4$. Let $\mathcal{P}_{1}$$\mathcal{P}_{2}$ be compatible partitions based at $v_{1}$$v_{2}$ respectively. Then we have a choice of sides $P_{1}^{\times}$$P_{2}^{\times}$ such that $P_{1}^{\times}\cap P_{2}^{\times}=\emptyset$. Observe that $P_{1}^{\times}$ cannot contain $C$, since $C\supset\{v_{2}^{\pm}\}$. Similarly, $P_{2}^{\times}$ must not contain the element $\{u_{1}^{\pm},v_{1}^{\pm},a_{1}^{\pm},b_{1}^{\pm}\}\in\mathcal{C}(v_{2})$. If we have $b_{2}\in P_{1}^{\times}$, then $P_{2}^{\times}$ must not contain $b_{2}$; the same holds for $\overline{b_{2}}$. Now, since in any compatible set, the partitions all based at $v_{1}$ must have their $v_{1}$-sides totally ordered by inclusion (and the same for $v_{2}$), we have $M(v_{1},v_{2})\leq 6$. Indeed, if we have four partitions based at $v_{1}$, then we may only add in a further two partitions based at $v_{2}$, both with one side only having components from $\{\{b_{3}\},\{\overline{b_{3}}\},\{b_{3}^{\pm}\}\}$ (for example, $\{v_{2},b_{3}\}$ and $\{v_{2},b_{3}^{\pm}\}$). The same holds if we are to have four partitions based at $v_{2}$. If we have three partitions based at $v_{1}$, then we can add in at most three partitions based at $v_{2}$: at least one of those based at $v_{1}$ necessarily contains at least one of $\{b_{2},\overline{b_{2}}\}$ on the same side as at least one of $\{b_{3},\overline{b_{3}}\}$. Hence $M(v_{1},v_{2})=6$.
@@
Consider any compatible set $\Pi$ of partitions based in the set $\{a_{1},a_{3}\}$. We consider how many partitions based at $a_{2}$ it is possible to add to $\Pi$ to obtain a set of size $M(a_{1},a_{2},a_{3})$. If $\Pi$ is empty, we can add at most five partitions based at $a_{2}$, since $M(a_{2})=5$. Assume $\Pi\neq\emptyset$; without loss of generality let $\mathcal{P}_{1}\in\Pi$ be based at $a_{1}$. We know that this has one side containing one or both of the components $\{u_{1}\}$$\{\overline{u_{1}}\}$. But these are both elements of $\mathcal{C}(a_{2})$, so if there is a partition based at $a_{1}$ with both of these components on the same side, then we can only add at most three partitions based at $a_{2}$. Similarly, if there is no partition based at $a_{1}$ with both of these components on the same side, then we may add at most four partitions based at $a_{2}$. Now a symmetric argument with any partitions based at $a_{3}$ further reduces the options for partitions based at $a_{2}$ which can be added to $\Pi$ to obtain a larger compatible set. If there is a partition based at $a_{3}$ with both of the components $\{u_{3}\}$$\{\overline{u_{3}}\}$ on the same side, then the number of partitions based at $a_{2}$ we can add is reduced by another two. If there is only one partition based at $a_{3}$ and the components $\{u_{3}\}$ and $\{\overline{u_{3}}\}$ are on different sides, then the number of partitions based at $a_{2}$ which we can add is further reduced by only one. Notice that if $\Pi$ contains two partitions based at $a_{1}$ then one of them necessarily has $\{u_{1}\}$ and $\{\overline{u_{1}}\}$ on the same side; the same holds for $a_{3}$. Hence $M(a_{1},a_{2},a_{3})\leq 5$ (and since $M(a_{2})=5$, in fact we have equality).
Hence $M(L)\leq M(v_{1},v_{2})+M(a_{1},a_{2},a_{3})=6+5=11$. We can attain this upper bound, as witnessed by the following compatible set of $\Delta$-partitions, defined by the sides:
@@
We would like to apply our work to show that $\textsc{vcd}(U(A_{\Delta}))=M(L)$. However, $\Delta$ violates Condition 2, so we cannot use Lemma 4.14. It turns out that $\Delta$ actually still satisfies the conclusion of Lemma 4.14; we now prove this by hand. Since our work only relies on graphs satisfying Condition 2 so that Lemma 4.14 is applicable, this is sufficient to apply Corollary 5.6.
Let $\Pi$ be a set of compatible $\Delta$-partitions. Suppose that $\Pi$ contains non-principal partitions $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ which are hugged in $\Pi$. We would like to show that if $\mathcal{R}$ is some principal partition which is compatible with all the principal partitions in $\Pi$ (so in particular, with all partitions involved in hugging $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$), then $\mathcal{R}$ is compatible with at least one of $\{\mathcal{Q},\mathcal{Q}^{\prime}\}$.